# Generalized Correlation Regression for Disentangling Dependence in Clustered Data

### v1

@@ Appendix D Simulations
D.1 Performance of the Proposed Approach
@@
with the logistic link function $\mu(x)=\frac{e^{x}}{1+e^{x}}$.
Under this hierarchical structure, the vector $\boldsymbol{\gamma}_{i}$ exhibits a block pattern:
[…]
so that $\alpha_{0}$ captures the baseline within-group correlation, while $\alpha_{1}$ represents the additional within-subgroup correlation. Consequently, the resulting correlation matrix ${\bf R}_{i}$ inherits the same block structure.
We set the true parameters to
@@
Summary across distributions. Taken together, the three simulation studies show that the proposed method performs reliably across Gaussian, Poisson, and Bernoulli responses. While binary data impose additional constraints and tend to exhibit slightly higher estimation error, the overall trends of decreasing MAE, stable coverage, and improved precision with larger $n$ are consistent with the theoretical asymptotics established in Section A.
D.2 Comparison with Competitors
@@
We consider binary outcomes with a logistic link function $\mu(x)=\frac{e^{x}}{1+e^{x}}$, so that $a^{\prime}(x)=\frac{e^{x}}{1+e^{x}}$ and $h(x)=x$. The group size is set as $m_{i}\sim\mathrm{Bin}(5,0.8)+2$, while the covariate vector ${\mathbf{x}}_{ij}$ and the linear predictor $\theta_{ij}$ follow the same setup as in Section D.1.
To ensure a fair comparison, unlike Section D.1, where the correlation matrices are generated under our proposed model, here we explicitly construct the correlation matrices under different pre-specified structures to mimic clustered and longitudinal data scenarios. This allows us to assess the robustness of each method when the true correlation deviates from the assumed working structure.
@@
Case 1: Two-level hierarchical clustering. In this setting, the data exhibit a two-level cluster structure similar to Section D.1. Each subject within a group belongs to one of two subgroups, indicated by
@@
Appendix E Additional real data examples
E.1 Toenail Infection Data
Variable Description
patientID Unique identifier for each patient in the trial
time The time in month when the visit actually took place: $[0,18.5]$
visit Number of visit attended: $\{1,2,3,4,5,6,7\}$
Table 11: Description of variables of the toenail infection dataset.
Parameter estimates for both models, along with their corresponding $p$-values, are reported in Table LABEL:table:data2_corr.
Variable Model (19) Model (20)
@@ E.1 Toenail Infection Data
treatment ($\beta_{1}$) $0.0313$ $0.8900$ $0.0292$ $0.9014$
time ($\beta_{2}$) 0.1530-0.1530*** $0.0000$ 0.1703-0.1703*** $0.0000$
For the correlation models, the results confirm a positive correlation between repeated measurements from the same patient, as captured by $\alpha_{0}$ or $\tilde{\alpha}_{0}$. This correlation decreases as the time gap between measurements increases, as reflected by the negative estimates of $\alpha_{2}$ or $\tilde{\alpha}_{2}$. The small estimates of $\alpha_{1}$ and $\tilde{\alpha}_{1}$ indicate that receiving the same treatment does not introduce additional correlation beyond the time effect, suggesting that the primary source of correlation is the patient-specific effect—an intuitively reasonable finding in this clinical context.
In this relatively standard setting, conventional methods such as the GEE and GLMM are directly applicable. We implement these approaches, using the gee and glmmTMB packages in R. For the GEE, we consider two working correlation structures: exchangeable and AR(1). To ensure stable estimation of the AR(1)-GEE model in gee, patients with only a single recorded visit are excluded from the analysis.
$\mathrm{logit}(p_{ijk})=\text{fixed effects}+u_{i}+v_{j}+w_{k},$
where $p_{ijk}=\mathbb{E}(y_{ijk})$ is the expected response for the $i$th patient, $j$th treatment type, and $k$th visit. Here, $u_{i}\sim\mathrm{N}(0,\sigma_{u}^{2})$ captures patient-level heterogeneity, $v_{j}\sim\mathrm{N}(0,\sigma_{v}^{2})$ captures treatment-level variation, and $w_{k}\sim\mathrm{N}(0,\sigma_{w}^{2})$ models visit-specific effects.
Fitting the models with both alternative approaches yields quantitatively similar results. We then compare their predictive performance against our proposed method. Predicted probabilities $\hat{p}_{ijk}$ for the GLMM are obtained using the predict function for glmmTMB. To assess predictive accuracy, we employ a 15-repeat stratified 5-fold cross-validation scheme. The dataset is first partitioned into two mutually exclusive subsets by treatment type, with random sampling conducted within each subset to preserve the original treatment distribution. Model performance is evaluated using the Brier Score and the Log Loss, with $\text{CV}_{1}$ and $\text{CV}_{2}$ denoting the mean cross-validated scores for the Brier Score and Log Loss, respectively; see Section 3 of the main part for their definitions.
Among all five models, our time-dependent correlation model (20) consistently delivers the best cross-validation performance, achieving the lowest mean CV scores with minimal variance for both metrics, while the GLMM performs worst. Pairwise $t$-tests comparing mean CV score differences between Model (20) and each of the other four models yield the $p$-values and $t$-statistics reported in Table 13. With the sole exception of the $\text{CV}_{1}$ comparison against the GEE (Exch) model, all differences are statistically significant at the 0.05 level. From an applied perspective, these results mean that our approach not only predicts patient outcomes more accurately but also clarifies how the timing of repeated measurements shapes correlation patterns over the treatment period. This combination of predictive strength and interpretability provides actionable insights for clinical trial design and patient monitoring strategies, offering a clear advantage over competing methods in both practical utility and statistical performance—likely attributable to its superior adequacy in model fitting.
Criteria Model (19) GEE (Exch) GEE (AR-1) GLMM
$\textbf{CV}_{1}$ 0.0004(-4.6182) 0.7990(-0.2595) 0.0003(-4.7687) 0.0000(-26.237)
$\textbf{CV}_{2}$ 0.0011(-4.1005) 0.0346(-2.3402) 0.0000(-6.5197) 0.0000(-31.865)
Table 13: Model comparison for toenail infection data.
E.2 Grouse Ticks Data
@@
The estimating results for our proposed method and the GLMM are reported in Table LABEL:table:data3_estimate, together with their corresponding p-values.
@@
From Tables LABEL:table:data3_estimate, all three models lead to consistent conclusions: (i) the covariates cHeight and Year have the same sign and significance pattern across models; and (ii) there is a significant clustering effect at the brood level, whereas the location-level effect is insignificant.

### v2

@@ Appendix C Simulations
C.1 Performance of the Proposed Approach
@@
with the logistic link function $\mu(x)=\frac{e^{x}}{1+e^{x}}$. We set the true parameters to
@@
Summary across distributions. Taken together, the three simulation studies show that the proposed method performs reliably across Gaussian, Poisson, and Bernoulli responses. While binary data impose additional constraints and tend to exhibit slightly higher estimation error, the overall trends of decreasing MAE, stable coverage, and improved precision with larger $n$ are consistent with the theoretical asymptotics established in Section B.
C.2 Comparison with Competitors
@@
We consider binary outcomes with a logistic link function $\mu(x)=\frac{e^{x}}{1+e^{x}}$, so that $a^{\prime}(x)=\frac{e^{x}}{1+e^{x}}$ and $h(x)=x$. The group size is set as $m_{i}\sim\mathrm{Bin}(5,0.8)+2$, while the covariate vector ${\mathbf{x}}_{ij}$ and the linear predictor $\theta_{ij}$ follow the same setup as in Section C.1.
To ensure a fair comparison, unlike Section C.1, where the correlation matrices are generated under our proposed model, here we explicitly construct the correlation matrices under different pre-specified structures to mimic clustered and longitudinal data scenarios. This allows us to assess the robustness of each method when the true correlation deviates from the assumed working structure.
@@
Case 1: Two-level hierarchical clustering. In this setting, the data exhibit a two-level cluster structure similar to Section C.1. Each subject within a group belongs to one of two subgroups, indicated by
@@
Appendix D Cross-validation Setup
We implement a 5-fold (stratified) cross-validation procedure repeated 15 times and compute two commonly used metrics: the Brier Score (Brier,, 1950) and the Log Loss, defined as
Brier Score $=\frac{1}{N}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}(y_{ij}-\hat{p}_{ij})^{2},$
where $n$ is the number of clusters, $m_{i}$ the size of cluster $i$, NN the total number of observations, $y_{ij}$ the observed outcome, and $\hat{p}_{ij}$ the predicted probability.
For each repetition, the performance is averaged over the 5 folds:
and the overall cross-validated score is computed as
$\text{CV}_{l}=\frac{1}{15}\sum_{r=1}^{15}\bar{s}_{r,l},\quad l=1,2,$
where $r$ indexes repetitions and $k$ indexes folds.
Appendix E Additional Analysis of Modern Prenatal Care Data
E.1 Analysis of the Full Correlation Models
We extend the initial baseline correlation model to incorporate socio-demographic and household-level covariates relevant to prenatal care choices:
$+\alpha_{12}\,w_{ijk,12}+\alpha_{13}\,w_{ijk,13},$ (23)
where $i$ index communities and $j,k$ index records,.
The additional covariates are defined as:
$+\alpha_{12}\,w_{ijk,12}+\alpha_{13}\,w_{ijk,13},$ (24)
where the notation follows that of (E.1).
The parameter estimates and corresponding $p$-values for models (E.1) and (E.1) are summarized in Table 14, facilitating comparison of the influence of generic versus covariate-specific correlation components.
Variable Model (E.1) Model (E.1)
@@ E.1 Analysis of the Full Correlation Models
Intercept ($\alpha_{0}$) $0.0471$. $0.0937$ 0.07390.0739* $0.0164$
mom ($\alpha_{1}$) 0.82660.8266*** $0.0000$
Ladino ($\alpha_{2}$) 0.0187-0.0187 $0.5628$ $0.0203$ $0.4893$
Unskilled ($\alpha_{5}$) $0.3592$ $0.3001$ 0.84540.8454*** $0.0000$
Professional ($\alpha_{6}$) $0.0756$ $0.7874$ 0.29300.2930** $0.0013$
Agri-self ($\alpha_{7}$) $0.0331$. $0.0823$ 0.04380.0438* $0.0102$
Agri-empl ($\alpha_{8}$) $0.0115$ $0.6601$ 0.05410.0541*** $0.0006$
Skilled ($\alpha_{9}$) $0.0629$ $0.3018$ 0.18420.1842*** $0.0000$
momEd ($\alpha_{10}$) $0.0066$ $0.6704$ 0.0279-0.0279* $0.0413$
husEd ($\alpha_{11}$) $0.0041$ $0.7744$ 0.0352-0.0352** $0.0023$
toilet ($\alpha_{12}$) $0.0073$ $0.8331$ 0.0413-0.0413 $0.2276$
TV ($\alpha_{13}$) 0.0004-0.0004 $0.9809$ 0.0186-0.0186 $0.2135$
Notes: . for $\text{p}<0.1$, * for $\text{p}<0.05$, ** for $\text{p}<0.01$ and *** for $\text{p}<0.001$.
Table 14: Parameter estimates and corresponding p-values for (E.1) and (E.1).
In Model (E.1), the hierarchical clustering effects $\hat{\alpha}_{0}$ and $\hat{\alpha}_{1}$ remain significant, while most additional family-level covariates are insignificant with near-zero effects, likely because the strong generic familial clustering $w_{ijk,1}$ captures most of their potential variation. Similarly, the two exceptions are $\hat{\alpha}_{4}$ (Indigenous Spanish ethnic-linguistic group) and $\hat{\alpha}_{7}$ (husband agri-self employment), both marginally significant (p <0.1<0.1), indicating additional dependence beyond generic family- and community-level effects.
For Model (E.1), which omits the generic familial clustering term $w_{ijk,1}$, several notable patterns emerge:
  1. 1.
    Effect of removing $w_{ijk,1}$: Without the generic familial clustering term, parameters for the family-level covariates generally have larger magnitudes and lower p-values compared to Model (E.1). This confirms that the model effectively partitions overall familial clustering into more detailed subcomponents.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
These results reveal insights often obscured by conventional models that capture only generic clustering. While community- and family-level dependence affects prenatal care choices, the underlying social and structural drivers typically remain hidden. Model (E.1) uncovers nuanced correlation patterns linked to social, ethnic, socio-demographic, and economic factors, highlighting mechanisms behind observed clustering and suggesting targets for health interventions or further investigation.
To examine individual-level factors that may influence correlations in prenatal care choices more closely, we extend Model (E.1) by incorporating pairwise differences in selected individual covariates:
where $i,j,k$ follow the same indexing as in (E.1). The additional covariates are defined as
$w_{ijk,16}=|\mathrm{birthOrd}_{ij}-\mathrm{birthOrd}_{ik}|,$
representing pairwise differences in maternal age, child age, and birth order, respectively. All other covariates remain as specified in (E.1). The corresponding parameter estimates and p-values are summarized in Table 15.
Variable Estimate P-value
Intercept ($\alpha_{0}$) 0.06830.0683* $0.0283$
mom ($\alpha_{1}$) 0.82690.8269*** $0.0000$
Ladino ($\alpha_{2}$) 0.0172-0.0172 $0.6092$
Indigenous non-Spanish ($\alpha_{3}$) 0.0034-0.0034 $0.9389$
Indigenous Spanish ($\alpha_{4}$) 0.0648-0.0648. $0.0792$
Unskilled ($\alpha_{5}$) $0.3768$ $0.2653$
Professional ($\alpha_{6}$) $0.0944$ $0.7185$
Agri-self ($\alpha_{7}$) $0.0293$ $0.1467$
Agri-empl ($\alpha_{8}$) $0.0166$ $0.5786$
Skilled ($\alpha_{9}$) $0.0578$ $0.3209$
momEd ($\alpha_{10}$) $0.0095$ $0.5445$
husEd ($\alpha_{11}$) $0.0027$ $0.8541$
toilet ($\alpha_{12}$) $0.0061$ $0.8722$
TV ($\alpha_{13}$) $0.0021$ $0.8981$
motherAge ($\alpha_{14}$) $0.0072$ $0.6238$
childAge ($\alpha_{15}$) 0.0251-0.0251 $0.1906$
birthOrd ($\alpha_{16}$) 0.0130-0.0130 $0.3412$
Notes: . for $\text{p}<0.1$, * for $\text{p}<0.05$, ** for $\text{p}<0.01$ and *** for $\text{p}<0.001$.
Table 15: Parameter estimates and corresponding p-values for (E.1).
Analysis of Model (E.1) yields results broadly consistent with our earlier findings. Although none of the additional individual-level covariates reach statistical significance, both $\hat{\alpha}_{15}$ and $\hat{\alpha}_{16}$ are estimated to be negative. This suggests a gradual weakening of shared prenatal care patterns over time and reduced similarity in care choices among children with larger age differences in larger families. Such patterns may reflect temporal changes in access, preferences, or health policies, as well as generational differences within multi-child households. However, given the limited data for these specific covariates, these observations should be interpreted cautiously, and further investigation is warranted before drawing definitive conclusions.
E.2 Comparisons with GLMM
We assess the performance of the GLMM as a main competing approach in this context. In addition to the baseline community- and family-level effects, we explore extensions incorporating additional hierarchical random effects. The GLMM is initially fitted using the lme4 package (Bates et al.,, 2015). For more complex specifications, where convergence issues arise due to model complexity, we employ the glmmTMB package, which is generally more robust for large models with multiple random effects. The corresponding results are presented in Tables 16 and 17.
Estimate Community Family indig husEmpl momEd husEd
Std.Dev. $0.0084$ $24.2176$ $0.0119$ $1.0357$ $0.0046$ $0.0020$
$p$-Value NA NA $0.4950$ NA NA NA
Table 16: Estimates of the random effects by the GLMM.
Estimate toilet TV motherAge childAge birthOrd
Std.Dev. $0.0127$ $0.0019$ $1.0726$ $3.3576$ $3.8209$
$p$-Value NA $1.0000$ $0.8731$ NA $1.0000$
Table 17: Estimates of the random effects by the GLMM.
Appendix F Analysis of Grouse Ticks Data
@@
The estimating results for our proposed method and the GLMM are reported in Table 19, together with their corresponding p-values.
@@
Notes: . for $\text{p}<0.1$, * for $\text{p}<0.05$, ** for $\text{p}<0.01$ and *** for $\text{p}<0.001$.
Table 19: Parameter estimates and corresponding p-values of our method and the GLMM.
From Tables 19, all three models lead to consistent conclusions: (i) the covariates cHeight and Year have the same sign and significance pattern across models; and (ii) there is a significant clustering effect at the brood level, whereas the location-level effect is insignificant.